The age of Jennilyn 3 years ago is one-fourth her age 6 years from now. How old is Jennilyn?
Answer by sachi(548)

The age of Jennilyn 3 years ago is one-fourth her age 6 years from now. How old is Jennilyn?
ans:-
let Jennilyn is x yrs old
3 years ago she was x-3 yrs
6 years from now she will be x+6 yrs
as per question x-3=(x+6)/4
or 4(x-3)=x+6
or 4x-12=x+6
or 3x=6+12=18
or x= 18/3=6 yrs ans